OpenStudy (anonymous):

subtract. simplify by collecting like radical terms if possible, assume that all number under radical represent non-negative number 3sqrt 81x - 3sqrt 3x^4 the 3 should be on top of the sqrt not in front of it

OpenStudy (anonymous):

Do I have it right?\[\sqrt[3]{81x}-\sqrt[3]{3^4}\]

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

3x^4

OpenStudy (anonymous):

3sqrt 3^4

OpenStudy (anonymous):

81=27*3 and 27 is a perfect cube (the cube root of 27 is 3 which will come out of the radical; also in the second term the x^4 = x^3*x, so\[3\sqrt[3]{3x}-x \sqrt[3]{3x}\](cont)

OpenStudy (anonymous):

you have missed the "x" in the second one

OpenStudy (anonymous):

These are known as like radical terms since they have the same index (3) and the same radicand 3x. That means we can simplify one step further by performing the subtraction\[(3-x)\sqrt[3]{3x}\]This is probably the form your teacher wants.

OpenStudy (anonymous):

I know I missed it in the begining, but I included it after.

OpenStudy (anonymous):

I got it thanks

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

I appreciate you help :-)

OpenStudy (anonymous):

anytime I'm on; that's why I'm here